### v1

@@ 5.3. How Effective Were AI-Generated Lesson Plans in the Classroom?
Teachers expressed high satisfaction with Shiksha copilot’s LPs: roughly 85% of plans were marked Very Good, 13% as Needs Some Improvement, and only 2% as Does Not Meet Requirements. However, as show in Figure 5, we observed notable differences across language mediums. English-medium plans received more Very Good ratings (87%) and fewer Needs Improvement ratings (11%) compared to Kannada-medium plans (81% and 18%, respectively). These differences were statistically significant ($\chi^{2}$ = 66.79, p ¡ 0.001, Cramer’s V = 0.110). Still, both mediums had very low rejection rates—roughly 3% for English and 1% for Kannada—indicating a high baseline quality. While these quantitative results point to broad satisfaction, interviews provided further insight into the usability, contextual alignment, and pedagogical value of the AI-generated content.
@@
Even though not all content in the LPs was directly applicable to every classroom, teachers appreciated the breadth and flexibility the plans offered. The inclusion of a wide range of examples, explanations, and activities allowed them to select those that best matched their teaching styles, student needs, and classroom contexts. Teachers valued having multiple options, which enabled them to tailor their instruction to different learning levels. This flexibility reduced the need to consult external resources and helped teachers deliver lessons more efficiently. Teacher 17 described:
@@
While AI-generated LPs featured diverse topics and pedagogical approaches, activity types—such as treasure hunts and matching exercises—frequently repeated across LPs. This repetition primarily stemmed from deliberate system design choices aimed at practicality in low-resource classrooms, where digital infrastructure and teaching materials are often limited or unavailable. Consequently, Shiksha copilot’s LPs were purposefully created to be usable and engaging without additional infrastructural or material demands (covered in more detail in Section 5.4.2). Although some curators perceived the repetition of activities as a limitation, teachers generally did not view it negatively; instead, they felt comfortable modifying activities and adding their own twist to better engage their students. Additionally, while the current system permitted teachers to edit only their individual copies of LPs, many expressed interest in sharing their classroom-tested activities to a collective resource pool, enriching available content and supporting the broader teaching community.
@@
Teachers appreciated that the LPs included content at varying levels of difficulty (beginner, intermediate, & advanced) which helped address the diverse learning needs within their classrooms. However, they also pointed out that these levels did not always align with their students’ actual competencies. In several instances, materials labeled as beginner still assumed foundational skills that many students had not yet developed. For example, a Grade 6 math teacher noted that the beginner-level materials required numeracy skills that their students were still acquiring. This mismatch reflects a broader limitation of AI-generated content: the system lacks access to real-time, classroom-specific learning data and therefore cannot fully adapt to the range of abilities present in different student groups. Teachers emphasized that even experienced educators often struggle to assess student levels without regular classroom interaction, underscoring the difficulty of pre-designing instructionally appropriate materials. According to Teacher 10, instructional planning is deeply situated, relational work that resists automation:
@@
Teachers valued that the LPs incorporated culturally and geographically relevant examples drawn from their own state. These references sparked student curiosity, made abstract ideas more relatable, and increased engagement. A science teacher shared how a science lesson explained the process of rusting through the example of coastal Karnataka, where salty air accelerates corrosion—making the concept both concrete and locally meaningful. Similarly, a Social Studies teacher appreciated how historical examples were connected to familiar places in Karnataka, helping students connect the material to their lived experiences. Teacher 11 highlighted how difficult it is to design LPs that reflect regional diversity and found the tool helpful in doing some of that work:
@@
At the same time, teachers expressed mixed views on the depth of contextualization. While state-level references were seen as valuable, some noted that examples from unfamiliar regions of Karnataka were not always meaningful to their students. Others felt the content lacked more immediate, hyper-local references tied to students’ neighborhoods or communities. A few teachers also suggested that the focus on state-level content came at the expense of national or global perspectives. Still, these concerns were seen relatively as minor. Most teachers felt confident in filling such contextual gaps themselves and emphasized that even partial regional grounding was a helpful starting point. These findings underscore that contextualization is not a one-time design task but an ongoing, situated process where teachers engage with their own expertise and AI outputs.
@@ 5.4. Broader Impact on Teachers’ Work Practices
The smartphone-based design of the tool introduced a level of flexibility that was previously unavailable. While teachers had earlier prepared LPs at home after school hours—often at the cost of personal or family time—they could now plan during commutes or in other spare moments throughout the work day. This shift allowed them to better manage their workload without encroaching on personal time. The convenience and support offered by the tool stood in stark contrast to previous educational interventions, which were often seen as adding to teachers’ burden. Teacher 8 with 20 years of teaching experience explained:
@@
Our survey results supported these qualitative findings. On average, teachers saved 2.02 hours per week on lesson planning ($\sigma=4.1$), with more active users saving 2.64 hours per week ($\sigma=4.0$) and less active users saving 1.50 hours per week ($\sigma=4.0$). These reductions showed a small to medium effect overall ($d=-0.494$, $r=0.336$), with a medium effect for active users ($d=-0.658$, $r=0.415$) and a small effect for less active users ($d=-0.371$, $r=0.262$).
Teachers also reported a measurable reduction in teaching-related stress. On a 7-point Likert scale, the average Teaching demands related stress score declined by 0.72 points ($\sigma=1.66$ ; d = -0.436; r = 0.289), indicating a small effect size. For example, teachers who previously responded Somewhat Agree to statements like ”I felt stressed for not having enough time to complete my teaching work (e.g., preparing, teaching the curricular content)” shifted closer to a Neutral response by the end of the intervention. As shown in the Figure  6, the post-intervention stress distribution not only shifted leftward but also became more concentrated in the lower stress range. Median scores decreased from 4.25 to 3.25. The effect was more pronounced among multi-day users, who showed an average decline of 0.85 points (d = -0.503), compared to a 0.60-point drop among single-day users (d = -0.371). These findings suggest that even modest reductions in stress were consistently experienced across user groups, with visible shifts away from moderate-to-high stress responses.
@@
The adoption of Shiksha copilot signaled a shift toward activity-based teaching. Although teachers already recognized the value of classroom activities for fostering student engagement and deeper learning, they often struggled to design and implement them consistently due to limited time and a lack of readily available resources. Shiksha copilot partially addressed this gap by providing easily accessible, contextually relevant activities tailored to each topic. These activities—such as visiting a gram panchayat office or collecting leaves with different venations—were grounded in the local environment and did not rely on digital infrastructure. Unlike device-dependent solutions such as Kahoot (Kah, [n. d.]) or Khanmigo (Kha, [n. d.]), the tool offered no-tech, practical ideas using commonly available materials, making them particularly well-suited for low-resource classrooms. One English teacher noted that the Clap and Tap activity—where students clap upon hearing consonant sounds and tap their heads for vowel sounds—was both engaging and required no materials. Similarly, physical activities like treasure hunts and relay races—where students search for concept-related clues around the school while competing with peers—supported embodied learning by combining cognitive and physical stimulation. Teachers emphasized that the availability of such ready-to-use activities reduced their lesson planning burden and motivated them to experiment with new instructional approaches. As one Kannada-medium teacher explained:
@@
While Shiksha copilot significantly reduced the time required for lesson planning, teachers emphasized that it had no impact on administrative tasks unrelated to teaching. Routine responsibilities like data entry and reporting continued to consume a large portion of their time. Even with easier access to high-quality, ready-to-use materials, many struggled to consistently implement activity-based teaching because administrative work often spilled into classroom hours—cutting into instructional time. One teacher pointed to the introduction of APAAR, a government initiative requiring detailed student ID data entry, as a recent example of added workload. In the absence of adequate clerical support, such tasks routinely fall to teachers. While the potential of Shiksha copilot to support more interactive and engaging pedagogy was widely recognized, teachers emphasized that its long-term success depends on systemic support—particularly reducing non-instructional burdens and improving staffing.
@@
The intervention also contributed to shifting teachers’ perceptions of AI, particularly around its ability to handle context-sensitive and local-language inputs. A senior teacher with 25 years of experience admitted that they initially associated AI with ”fake videos and images,” but after using Shiksha copilot, came to see it as a practical and accessible tool for quickly retrieving relevant information. There was initial skepticism about whether AI could generate content that resonated with students, especially in regional languages like Kannada. One teacher shared that they had assumed such technologies would not function well in Kannada but were pleasantly surprised by Shiksha copilot’s ability to interpret and respond appropriately to local-language inputs. These experiences suggest that while baseline trust in AI was relatively high, hands-on interaction with Shiksha copilot helped demystify its capabilities and enhance its perceived usefulness in everyday teaching. At the same time, teachers expressed concerns about overreliance. One experienced teacher joked that she worried she might “stop thinking about the topics” because the AI provided information so quickly and easily. Others emphasized the importance of learning to use AI tools thoughtfully—not only to stay current alongside their students but also to exercise professional judgment to ensure these technologies are used in pedagogically beneficial ways.
@@ 6.2. Leveraging the Communities of Practice
The concept of Communities of Practice (CoPs), introduced by Lave and Wenger (Lave, 1991), offers a valuable lens for understanding how teaching is socially constructed through shared activitiy and meaning-making. CoPs emerge through mutual engagement, a joint enterprise, and a shared repertoire of practices and resources developed through sustained participation (Wenger, 1999). In education, CoPs have been shown to support pedagogical innovation, deepen teacher expertise, and foster sustained teacher learning and growth (Supovitz, 2002; Paltiwale et al., 2020).
@@ 7. Conclusion
This study examined how teachers in low-resource classrooms integrate an AI-based lesson planning tool into their instructional practices. By deploying the tool with over 1,000 teachers, we gained valuable insights into how AI-generated LPs transform teaching workflows in challenging environments. Our findings reveal that the AI system significantly reduced administrative burden while simultaneously enhancing instructional quality. Teachers strategically leveraged the curriculum-aligned content as a foundation, enriching it with localized examples, interactive activities, and tailored assessments to meet their students’ specific needs. The tool’s mobile-optimized, low-tech design proved particularly effective within infrastructural constraints of low-resource schools, allowing teachers to create more engaging lessons with reduced preparation time. The study also identified important limitations. Language support for vernacular content presented challenges, with Kannada LPs requiring substantial modifications compared to English materials. Additionally, while individual teachers benefited from the tool, its impact was constrained by broader systemic issues including staffing shortages and administrative pressures. The largely individual usage pattern, without sustained mentoring or collaborative frameworks, limited the tool’s potential for deeper pedagogical transformation. These findings highlight that effective AI integration in education requires more than technological solutions. Success depends on embedding these tools within existing communities of practice where teachers can collectively adapt, reflect upon, and share knowledge about AI-assisted teaching. By positioning AI as a participatory resource aligned with teachers’ values and expertise rather than a standalone solution, we can better support educational improvement in low-resource contexts.
Acknowledgements
We would like to thank the teachers who participated in the study for their time and thoughtful inputs. We extend our acknowledgments to Sikshana Foundation, especially Srinivas Mysore and Rijutha Raj, for their implementation support.

### v2

@@ 5.3. How Effective Were AI-Generated Lesson Plans in the Classroom?
Teachers expressed high satisfaction with Shiksha Copilot’s LPs: roughly 85% of plans were marked Very Good, 13% as Needs Some Improvement, and only 2% as Does Not Meet Requirements. However, as shown in Figure 7, we observed notable differences across language mediums. English-medium plans received more Very Good ratings (87%) and fewer Needs Improvement ratings (11%) compared to Kannada-medium plans (81% and 18%, respectively). These differences were statistically significant ($\chi^{2}$ = 66.79, p ¡ 0.001, Cramer’s V = 0.110). Still, both mediums had very low rejection rates—roughly 3% for English and 1% for Kannada—indicating a high baseline quality. While these quantitative results point to broad satisfaction, interviews provided further insight into the usability, contextual alignment, and pedagogical value of the AI-generated content.
@@
Even though not all content in the LPs was directly applicable to every classroom, teachers appreciated the breadth and flexibility the plans offered. The inclusion of a wide range of examples, explanations, and activities allowed them to select those that best matched their teaching styles, student needs, and classroom contexts. Teachers valued having multiple options, which enabled them to tailor their instruction to different learning levels. This flexibility reduced the need to consult external resources and helped teachers deliver lessons more efficiently. Teacher 17, an English teacher with 21 years of teaching experience, described:
@@
While AI-generated LPs featured diverse topics and pedagogical approaches, activity types—such as treasure hunts and matching exercises—frequently repeated across LPs. This repetition primarily stemmed from deliberate system design choices aimed at practicality in low-resource classrooms, where digital infrastructure and teaching materials are often limited or unavailable. Consequently, Shiksha Copilot’s LPs were purposefully created to be usable and engaging without additional infrastructural or material demands (covered in more detail in §5.4.2). Although some curators perceived the repetition of activities as a limitation, teachers generally did not view it negatively; instead, they felt comfortable modifying activities and adding their own twist to better engage their students. Additionally, while the current system permitted teachers to edit only their individual copies of LPs, many expressed interest in sharing their classroom-tested activities to a collective resource pool, enriching available content and supporting the broader teaching community.
@@
Teachers appreciated that the LPs included content at varying levels of difficulty (beginner, intermediate, & advanced) which helped address the diverse learning needs within their classrooms. However, they also pointed out that these levels did not always align with their students’ actual competencies. In several instances, materials labeled as beginner still assumed foundational skills that many students had not yet developed. For example, a Grade 6 math teacher noted that the beginner-level materials required numeracy skills that their students were still acquiring. This mismatch reflects a broader limitation of AI-generated content: the system lacks access to real-time, classroom-specific learning data and therefore cannot fully adapt to the range of abilities present in different student groups. Teachers emphasized that even experienced educators often struggle to assess student levels without regular classroom interaction, underscoring the difficulty of pre-designing instructionally appropriate materials. According to Teacher 10, a Mathematics teacher, lesson planning is deeply situated, relational work that resists automation:
@@
Teachers valued that the LPs incorporated culturally and geographically relevant examples drawn from their own state. These references sparked student curiosity, made abstract ideas more relatable, and increased engagement. A science teacher shared how a science lesson explained the process of rusting through the example of coastal Karnataka, where salty air accelerates corrosion—making the concept both concrete and locally meaningful. Similarly, a Social Studies teacher appreciated how historical examples were connected to familiar places in Karnataka, helping students connect the material to their lived experiences. Teacher 11, a Social Studies teacher, highlighted how difficult it is to design LPs that reflect regional diversity and found the tool helpful in doing some of that work:
@@
Difficulties with local contextualization. At the same time, teachers expressed mixed views on the depth of contextualization. While state-level references were seen as valuable, some noted that examples from unfamiliar regions of Karnataka were not always meaningful to their students. Others felt the content lacked more immediate, hyper-local references tied to students’ neighborhoods or communities. This limitation in achieving hyper-local contextualization stemmed from the design of the RAG system, which generated content intended to cover the entire state at scale. Although teachers had both UI- and AI-based affordances to update content for their local context, they more often made mental adaptations off-platform to suit their students’ needs, as discussed in §5.1.3. A few teachers also suggested that the focus on state-level content came at the expense of national or global perspectives. Still, these concerns were seen relatively as minor. Most teachers felt confident in filling such contextual gaps themselves and emphasized that even partial regional grounding was a helpful starting point. These findings underscore that contextualization is not a one-time design task but an ongoing, situated process where teachers engage with their own expertise and AI outputs.
@@ 5.4. Broader Impact on Teachers’ Work Practices
Increased convenience. The smartphone-based design of the tool introduced a level of flexibility that was previously unavailable. While teachers had earlier prepared LPs at home after school hours—often at the cost of personal or family time—they could now plan during commutes or in other spare moments throughout the work day. This shift allowed them to better manage their workload without encroaching on personal time. The convenience and support offered by the tool stood in stark contrast to previous educational interventions, which were often seen as adding to teachers’ burden. Teacher 8, a Science teacher, with 20 years of teaching experience explained:
@@
Our survey results supported these qualitative findings. On average, teachers saved 2.02 hours per week on lesson planning ($\sigma=4.1$), with more active users saving 2.64 hours per week ($\sigma=4.0$) and less active users saving 1.50 hours per week ($\sigma=4.0$). These reductions showed a small to medium effect overall (d=0.494d=-0.494, $r=0.336$), with a medium effect for active users (d=0.658d=-0.658, $r=0.415$) and a small effect for less active users (d=0.371d=-0.371, $r=0.262$).
Reduced stress. Teachers also reported a measurable reduction in teaching-related stress. On a 7-point Likert scale, the average Teaching demands related stress score declined by 0.72 points ($\sigma=1.66$ ; d = -0.436; r = 0.289), indicating a small effect size. For example, teachers who previously responded Somewhat Agree to statements like ”I felt stressed for not having enough time to complete my teaching work (e.g., preparing, teaching the curricular content)” shifted closer to a Neutral response by the end of the intervention. As shown in the Figure  8, the post-intervention stress distribution not only shifted leftward but also became more concentrated in the lower stress range. Median scores decreased from 4.25 to 3.25. The effect was more pronounced among multi-day users, who showed an average decline of 0.85 points (d = -0.503), compared to a 0.60-point drop among single-day users (d = -0.371). These findings suggest that even modest reductions in stress were consistently experienced across user groups, with visible shifts away from moderate-to-high stress responses.
@@
The adoption of Shiksha Copilot signaled a shift toward activity-based teaching. Teachers had been using the 5E lesson plan format even before the introduction of Shiksha Copilot, and they already recognized the value of classroom activities for fostering student engagement and deeper learning. However, they often struggled to design and implement such activities consistently due to limited time and a lack of readily available resources. Shiksha Copilot partially addressed this gap by providing easily accessible, contextually relevant activities tailored to each topic. These activities—such as visiting a gram panchayat office or collecting leaves with different venations—were grounded in the local environment and did not rely on digital infrastructure. Unlike device-dependent solutions such as Kahoot (Kah, [n. d.]) or Khanmigo (Kha, [n. d.]), the tool offered no-tech, practical ideas using commonly available materials, making them particularly well-suited for low-resource classrooms. One English teacher noted that the Clap and Tap activity—where students clap upon hearing consonant sounds and tap their heads for vowel sounds—was both engaging and required no materials. Similarly, physical activities like treasure hunts and relay races—where students search for concept-related clues around the school while competing with peers—supported embodied learning by combining cognitive and physical stimulation. Teachers emphasized that the availability of such ready-to-use activities reduced their lesson planning burden and motivated them to experiment with new instructional approaches. Teacher 15, a Social Science teacher, explained:
@@
Contingent on systemic factors. While Shiksha Copilot significantly reduced the time required for lesson planning, teachers emphasized that it had no impact on administrative tasks unrelated to teaching. Routine responsibilities like data entry and reporting continued to consume a large portion of their time. Even with easier access to high-quality, ready-to-use materials, many struggled to consistently implement activity-based teaching because administrative work often spilled into classroom hours—cutting into instructional time. One teacher pointed to the introduction of APAAR, a government initiative requiring detailed student ID data entry, as a recent example of added workload. In the absence of adequate clerical support, such tasks routinely fall to teachers. While the potential of Shiksha Copilot to support more interactive and engaging pedagogy was widely recognized, teachers emphasized that its long-term success depends on systemic support—particularly reducing non-instructional burdens and improving staffing. Teacher 14, a Science and Mathematics teacher, succinctly put it:
There is no challenge for teachers in implementing these LPs. Teacher shortage is the main problem. Now we have only three teachers. One of our H.M. [Headmaster] went for training. Only three teachers, we handle seven classes. That is a government problem.
@@
Shift in perception of AI. The intervention also contributed to shifting teachers’ perceptions of AI, particularly around its ability to handle context-sensitive and local-language inputs. A senior teacher with 25 years of experience admitted that they initially associated AI with ”fake videos and images,” but after using Shiksha Copilot, came to see it as a practical and accessible tool for quickly retrieving relevant information. There was initial skepticism about whether AI could generate content that resonated with students, especially in regional languages like Kannada. One teacher shared that they had assumed such technologies would not function well in Kannada but were pleasantly surprised by Shiksha Copilot’s ability to interpret and respond appropriately to local-language inputs. These experiences suggest that while baseline trust in AI was relatively high, hands-on interaction with Shiksha Copilot helped demystify its capabilities and enhance its perceived usefulness in everyday teaching. At the same time, teachers expressed concerns about overreliance. One experienced teacher joked that she worried she might “stop thinking about the topics” because the AI provided information so quickly and easily. Others emphasized the importance of learning to use AI tools thoughtfully—not only to stay current alongside their students but also to exercise professional judgment to ensure these technologies are used in pedagogically beneficial ways.
@@ 6.2. Leveraging the Communities of Practice
The concept of Communities of Practice (CoPs), introduced by Lave and Wenger (Lave, 1991), offers a valuable lens for understanding how teaching is socially constructed through shared activity and meaning-making. CoPs emerge through mutual engagement, a joint enterprise, and a shared repertoire of practices and resources developed through sustained participation (Wenger, 1999). In education, CoPs have been shown to support pedagogical innovation, deepen teacher expertise, and foster sustained teacher learning and growth (Supovitz, 2002; Paltiwale et al., 2020).
@@ 7. Conclusion
This study examined how teachers in low-resource classrooms integrate an AI-based lesson planning tool into their instructional practices. By deploying the tool with over 1,000 teachers, we gained valuable insights into how AI-generated LPs transform teaching workflows in challenging environments. Our findings reveal that the AI system significantly reduced administrative burden while simultaneously enhancing instructional quality. Teachers strategically leveraged the curriculum-aligned content as a foundation, enriching it with localized examples, interactive activities, and tailored assessments to meet their students’ specific needs. The tool’s mobile-optimized, low-tech design proved particularly effective within infrastructural constraints of low-resource schools, allowing teachers to create more engaging lessons with reduced preparation time. The study also identified important limitations. Language support for vernacular content presented challenges, with Kannada LPs requiring substantial modifications compared to English materials. Additionally, while individual teachers benefited from the tool, its impact was constrained by broader systemic issues including staffing shortages and administrative pressures. The largely individual usage pattern, without sustained mentoring or collaborative frameworks, limited the tool’s potential for deeper pedagogical transformation. These findings highlight that effective AI integration in education requires more than technological solutions. Success depends on embedding these tools within existing communities of practice where teachers can collectively adapt, reflect upon, and share knowledge about AI-assisted teaching. By positioning AI as a participatory resource aligned with teachers’ values and expertise rather than a standalone solution, we can better support educational improvement in low-resource contexts. The tool has continued to evolve beyond the study period, with ongoing refinements shaped by the insights we gathered. It has also continued to scale to reach more teachers as we explore pathways to broaden its use in ways that remain responsive to their needs and contextual realities.